# A Flat Minima Perspective on Understanding Augmentations and Model Robustness

### v1

@@ 8 Acknowledgment
This work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2020-II201336, Artificial Intelligence Graduate School Program (UNIST)), (No. IITP-2025-RS-2022-00156361, Innovative Human Resource Development for Local Intellectualization program), and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00459023).
Appendix A Supplementary Materials
@@
This appendix contains additional material that could not be incorporated into the main paper due to page constraints, including detailed proofs of the theorems and experimental details. Section B provides detailed proofs of Theorems 1 to 3, Section C presents experimental details and additional results, Section D addresses related works in detail, and Section E offers further discussion on broader applicability of our findings including practical usage.
Appendix B Proofs on Theorems
B.1 Proof on Theorem 1
@@
Using the first-order Taylor expansion, $f(x;\theta+\Delta)\approx f(x;\theta)+\mathbf{J}_{\boldsymbol{\theta}}\cdot\Delta$. This approximation is supported by two main observations. First, when $f$ is a linear model, the Taylor expansion becomes exact, and a linear model with infinite width serves as a universal approximator. Second, supported by lazy training regime [8, 13, 35, 3] and Neural Tangent Kernel (NTK) [25, 32] statements, training loss for a deep model rapidly converges to zero with minimal parameter updates. The neural network is known to behave like its linearization around the initial weight in the entire training process, and the change in Jacobian becomes negligible.
@@
By Sard’s theorem [45]—which states that the set of critical values of a smooth function has Lebesgue measure zero—the Jacobian matrices have full rank almost everywhere. Therefore, we will assume the matrices $\mathbf{J}_{\mathbf{x}}$ and $\mathbf{J}_{\boldsymbol{\theta}}$ to be full-rank matrices as in [58, 14]. Since $c\ll n,p$ and $\mathbf{J}_{\boldsymbol{\theta}},\mathbf{J_{x}}$ are full-rank, there always exists a solution $\delta$ for the equation (6). Using psuedoinverse notation $\mathbf{J}_{\mathbf{x}}^{+}$ and kernel notation $\text{ker}(\mathbf{J_{x}})$ for $\mathbf{J_{x}}$, the solution for equation (6) can be written as
@@
We employ the same strategy to that in the proof of Theorem 1.A. (Appendix B.1.1) For conciseness, we do not restate the complete set of assumptions and justifications presented previously. Under the first-order Talyor expansion, $f(x;\theta+\Delta)\approx f(x;\theta)+\mathbf{J}_{\boldsymbol{\theta}}\cdot\Delta$ and $f(x+\delta;\theta)\approx f(x;\theta)+\mathbf{J_{x}}\cdot\delta$, resulting in $\mathbf{J}_{\boldsymbol{\theta}}\cdot\Delta=\mathbf{J_{x}}\cdot\delta$. Since $\mathbf{J}_{\boldsymbol{\theta}}$ and $\mathbf{J_{x}}$ are full-rank, the solution for the equation becomes
@@ B.2 Proof on Theorem 2
Let $\theta^{*}_{0}$ be an optimal parameter of a model that satisfies for all $(x,y)\in\mathcal{D},$ $\mathcal{L}(f(x;\theta_{0}^{*}),y)$ $=0$ and for all $(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{c}\ /\ \mathcal{D},$ $\mathcal{L}(f(x;\theta_{0}^{*}),y)>0$. Let $b_{0}^{*}$ denote the $b$-value of the $b$-flat minima for $\theta^{*}_{0}$.
@@
[…]
@@
as follows. By definition, $\gamma_{\mathcal{A}}$ represents the maximum $L_{2}$ distance that augmentation has non-zero probabilities around the original input. Define $\mathcal{C}_{\mathcal{X}}=\{\delta\in\mathbb{R}^{n}:\|\delta\|\leq\gamma_{\mathcal{A}}\}.$ Under Theorem 1, $\mathcal{C}_{\mathcal{X}}$ is the compensatory set for the parameter perturbations
@@ B.3 Proof on Theorem 3
(Generalization bound) Given $M$ covering sets $\{\Theta_{k}\}_{k=1}^{M}$ of parameter space $\Theta$ with $\Theta=\bigcup_{k=1}^{M}\Theta_{k}$ and $\text{diam}(\Theta)=\sup_{\theta,\theta^{\prime}\in\Theta}||\theta-\theta^{\prime}||_{2}$, where $M=\left\lceil\frac{\text{diam}(\Theta)}{\gamma_{\Theta}}\right\rceil^{d}$, and VC dimension $v_{k}$ for each $\Theta_{k}$, the following inequality holds with probability at least $1-\delta$:
$\mathcal{E}_{\mathcal{T}}(\tilde{\theta}^{*})$ <^$<\hat{\mathcal{E}}_{\mathcal{\tilde{D}}}(\tilde{\theta}^{*})+\frac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})+\max_{k\in[1,M]}\left[\sqrt{\frac{v_{k}\log(N/v_{k})}{N}+\frac{\log(M/\delta)}{N}}\right],$
@@
Remind that $\mathcal{\tilde{D}}$ represents augmented data distribution, and $\gamma_{\Theta}:=\left(\max_{x\in\mathcal{D}}\sigma_{\max}^{\theta}/\sigma_{\min}^{x}\right)^{-1}\cdot\gamma_{\mathcal{A}}$ the radius of the compensatory set $\mathcal{C}_{\Theta}$ defined in Theorem 1. The robust empirical risk is originally defined [7] as:
^$\hat{\mathcal{E}}_{\mathcal{D}}^{\gamma}(\theta):=\max_{\|\delta\|\leq\gamma}\frac{1}{N}\sum_{i=1}^{N}\big{[}L(f(x_{i},\theta+\delta),y_{i})\big{]}.$
@@
^$\hat{\mathscr{E}}_{\mathcal{D}}^{\gamma}(\theta):=\max_{\|\delta\|\leq\gamma}\frac{1}{N}\sum_{i=1}^{N}\big{[}L(f(x_{i}+\delta,\theta),y_{i})\big{]}.$
@@
We now list the necessary propositions from [7] for the rest of the proof.
Proposition 1. For all $\theta\in\Theta,\ |\mathcal{E}_{\mathcal{D}}(\theta)-\mathcal{E}_{\mathcal{T}}(\theta)|\leq\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})$
@@
$\mathcal{E}_{\mathcal{T}}(\tilde{\theta}^{*})$ $\leq\mathcal{E}_{\mathcal{D}}(\tilde{\theta}^{*})+\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})$ Prep. 1
@@
^$\leq\hat{\mathscr{E}}_{\mathcal{D}}^{\gamma_{\mathcal{A}}}(\tilde{\theta}^{*})+\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})+\mathcal{B}(M,N,\delta)$ Def. of $\mathscr{E},\gamma_{\Theta},$ Thm. 1
^$\leq\hat{\mathcal{E}}_{\tilde{\mathcal{D}}}(\tilde{\theta}^{*})+\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})+\mathcal{B}(M,N,\delta)$ Def. of $\gamma_{\mathcal{A}},\,\tilde{\theta}^{*}$
Appendix C Experimental Details and Additional Experiments
In this section, we provide a comprehensive account of the experiments corresponding to the tables and figures presented in the main paper, along with supplementary results and discussions that were excluded from the primary manuscript due to page constraints. The additional results include extra robustness evaluations on different backbone networks, flatness evaluations on (tiny)ImageNet belonging to Table 2, clean accuracies of models in Table 3, and singular value bounds on Theorem 1.
C.1 Detailed Experimental Methodology
For all the experiments assessing flatness, mean accuracy, mean corruption error (mCE), and adversarial robustness, we trained three independent models per benchmark and reported the averaged results, with the exception of the ImageNet experiment in Table 3, which was conducted only once.
C.1.1 Flatness Metrics
@@
Training Details. Our training procedure closely follows [21], with the only difference being that we train the WideResNet-40-2 model for 200 epochs instead of 100. We employ an SGD optimizer with an initial learning rate of 0.1, a weight decay of 0.0005, and a momentum of 0.9. Additionally, we use a cosine learning rate decay scheduler, starting with a learning rate of 0.1 and a weight decay of 0.0005. Finally, we evaluate and report the flatness of the converged models.
C.1.2 Robustness Benchmark Details
Domain Generalization Benchmarks. We conducted comprehensive domain generalization experiments on the PACS and OfficeHome datasets to evaluate model robustness and generalizability using augmentations. These datasets are specifically designed to test a model’s ability to generalize across different domains, as they consist of images that share the same labels but differ in their domain representations. For example, the PACS dataset contains images from four distinct domains—’art painting’, ’cartoon’, ’photo’, and ’sketch’—all featuring the same classes such as ’dog’, ’elephant’, and others. In all experiments, models were trained until convergence to ensure fair and consistent comparisons across different settings.
Our baseline employed Empirical Risk Minimization (ERM) with the best hyperparameters selected from DomainBed’s ERM training configuration [17]. To specifically assess the impact of augmentations on model performance, we excluded the heavy augmentation compositions used in the original DomainBed configuration.
We have adopted 2,500 steps of training iterations, which results in slightly faster training than the default setting of 5,000. Specifically for StyleAug, we performed additional tuning of its hyperparameter $\alpha$ alongside the training epochs. Through manual experimentation, we found that setting $\alpha=1.0$ yielded the best results in terms of model robustness and generalization performance.
Common Corruption Benchmarks. We assessed the common corruption robustness of models trained with augmentations on the CIFAR-10/100, tinyImageNet, and ImageNet datasets. Specifically, we utilized their corrupted counterparts—CIFAR-10/100-C/$\overline{\text{C}}$, tinyImageNet-C, and ImageNet-C— which are standard benchmarks generated from the original datasets by applying 15 distinct corruption types at 5 different severity levels. These corruption types include brightness changes, contrast alterations, defocus blur, elastic transformations, fog addition, frost addition, Gaussian blur, glass distortion, impulse noise, JPEG compression, motion blur, pixelation, shot noise, snow addition, and zoom blur. Models were trained and validated on the respective clean datasets, with robustness evaluations performed at test time only to simulate real-world scenarios where models encounter unforeseen corruptions. To quantify model performance across all corruptions and severity levels, we calculated the mean Corruption Error (mCE), which averages the error rates over the different conditions.
For CIFAR-10/100 datasets, the training details are exactly the same as in C.1.1. For tinyImageNet and ImageNet experiments, we have majorly adopted the configurations from [21], with changes in epochs and backbone networks for faster training. We have trained ResNet18 on tinyImageNet and ImageNet for 50 epochs, instead of training ResNet50 for 90 epochs. SGD optimizer with initial learning rate 0.01, weight decay 0.0005, and momentum 0.9 has been adopted. All the models have been trained until convergence, with cosine learning rate decaying scheduling.
Regarding the augmentation techniques, default configuration parameters were used for all baseline augmentations except StyleAug, which had hyperparameter $\alpha$ set to $0.95$ for the CIFAR experiments, and 1.0 for the tinyImageNet experiments.
@@ C.1.2 Robustness Benchmark Details
Adversarial Robustness Benchmarks. To evaluate the robustness of models in adversarial settings, we trained models with each augmentation on CIFAR-10/100, tinyImageNet, and ImageNet, and conducted tests using untargeted Projected Gradient Descent (PGD) attacks based on the $L_{2}$ and $L_{\infty}$ norm. The training environment holds the same as in the common corruption benchmarks. We will use $\epsilon$ and $\alpha$ to denote the maximum allowed perturbation size.
Regarding CIFAR-10/100 experiments, we utilized attack configurations from [55]. Specifically, we deployed PGD-20 $L_{2}$ attacks with $\epsilon=0.5,\alpha=0.0125$ and PGD-7 $L_{\infty}$ attacks with $\epsilon=8/255,\alpha=2/255$. For tinyImageNet experiments, we utilized PGD-10 $L_{2}$ attacks with $\epsilon=0.25,\alpha=0.025$ and PGD-5 $L_{\infty}$ attacks with $\epsilon=8/255,\alpha=2/255$. On ImageNet experiments, we deployed PGD-10 $L_{2}$ attacks with $\epsilon=0.25,\alpha=0.025$ and PGD-2 $L_{\infty}$ attacks with $\epsilon=2/255,\alpha=1/255$. In essence, milder attack configurations have been deployed for tinyImageNet and ImageNet.
C.1.3 Results on Theorem 1’s Bound

### v3

@@ Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00459023), Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2020-II201336, Artificial Intelligence Graduate School Program at UNIST), (No. RS-2025-25442824, AI Star Fellowship Program at UNIST), (No. RS-2025-25441996, Development of a Virtual Tactile Signal Generation Platform Technology Based on Multimodal Vision-Tactile Integrated AGI), (No. IITP-2025-RS-2022-00156361, Innovative Human Resource Development for Local Intellectualization program).
A. Supplementary Materials
@@
This appendix contains additional material that could not be incorporated into the main paper due to page constraints, including detailed proofs of the theorems and experimental details. Section B provides detailed proofs of Theorems 1 to 3, Section C presents experimental details and additional results, and Section D offers further discussion on broader applicability of our findings including practical usage.
B. Proofs on Theorems
B.1. Proof on Theorem 1
@@
Using the first-order Taylor expansion, $f(x;\theta+\Delta)\approx f(x;\theta)+\mathbf{J}_{\boldsymbol{\theta}}\cdot\Delta$. This approximation is supported by two main observations. First, when $f$ is a linear model, the Taylor expansion becomes exact, and a linear model with infinite width serves as a universal approximator. Second, supported by lazy training regime (lazytraining0; lazytraining1; lazytraining2; lazytraining4) and Neural Tangent Kernel (NTK) (NTK0; NTK1) statements, training loss for a deep model rapidly converges to zero with minimal parameter updates. The neural network is known to behave like its linearization around the initial weight in the entire training process, and the change in Jacobian becomes negligible.
@@
A neural network with standard deep learning activations (e.g. ReLU family, sigmoid/tanh family, softplus, GELU, etc.) is said to be in o-minimal structure (or tame)  (Tame1; Tame2), which enables us to utilize Sard’s theorem. By Sard’s theorem (Sard; MorseSard) stating that the set of critical values has Lebesgue measure zero, the Jacobian matrices have full rank almost everywhere. Therefore, we will assume the matrices $\mathbf{J}_{\mathbf{x}}$ and $\mathbf{J}_{\boldsymbol{\theta}}$ to be full-rank matrices as in (jacobianfullrank0; jacobianfullrank1). Since $c\ll n,p$ and $\mathbf{J}_{\boldsymbol{\theta}},\mathbf{J_{x}}$ are full-rank, there always exists a solution $\delta$ for the equation (6). Using psuedoinverse notation $\mathbf{J}_{\mathbf{x}}^{+}$ and kernel notation $\text{ker}(\mathbf{J_{x}})$ for $\mathbf{J_{x}}$, the solution for equation (6) can be written as
@@
We employ the same strategy to that in the proof of Theorem 1.A. For conciseness, we do not restate the complete set of assumptions and justifications presented previously. Under the first-order Talyor expansion, $f(x;\theta+\Delta)\approx f(x;\theta)+\mathbf{J}_{\boldsymbol{\theta}}\cdot\Delta$ and $f(x+\delta;\theta)\approx f(x;\theta)+\mathbf{J_{x}}\cdot\delta$, resulting in $\mathbf{J}_{\boldsymbol{\theta}}\cdot\Delta=\mathbf{J_{x}}\cdot\delta$. Since $\mathbf{J}_{\boldsymbol{\theta}}$ and $\mathbf{J_{x}}$ are full-rank, the solution for the equation becomes
@@ B.2. Proof on Theorem 2
Let $\theta^{*}_{0}$ be an optimal parameter of a model that satisfies for all $(x,y)\in\mathcal{D},$ $\mathcal{L}(f(x;\theta_{0}^{*}),y)$ =0=0 and for all $(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{c}\ /\ \mathcal{D},$ $\mathcal{L}(f(x;\theta_{0}^{*}),y)>0$. Let $b_{0}^{*}$ denote the $b$-value of the $b$-flat minima for $\theta^{*}_{0}$.
@@
, where $\lambda>0$. Then, for all $\|\delta\|\leq\lambda\cdot b_{0}^{*},\ \mathcal{L}(f(x+\delta;\theta_{0}^{*}),y)=\mathcal{L}(f(x;\theta_{0}^{*}),y))=0$ by the definition of functional compensatory set. Nonetheless, $\mathcal{L}(f(x+\delta;\theta_{0}^{*}),y)>0$ by the definition of $\theta_{0}^{*}$. $\Rightarrow\Leftarrow$
@@
as follows. By definition, $\gamma_{\mathcal{A}}$ represents the maximum $L_{2}$ distance that augmentation has nonzero probabilities around the original input. Define $\mathcal{C}_{\mathcal{X}}=\{\delta\in\mathbb{R}^{n}:\|\delta\|\leq\gamma_{\mathcal{A}}\}.$ Under Theorem 1, $\mathcal{C}_{\mathcal{X}}$ is the compensatory set for the parameter perturbations
@@ B.3. Proof on Theorem 3
[…]
@@
Remind that $\mathcal{\tilde{D}}$ represents augmented data distribution, and $\gamma_{\Theta}:=\left(\max_{x\in\mathcal{D}}\sigma_{\max}^{\theta}/\sigma_{\min}^{x}\right)^{-1}\cdot\gamma_{\mathcal{A}}$ the radius of the compensatory set $\mathcal{C}_{\Theta}$ defined in Theorem 1. The robust empirical risk is originally defined (SWAD:2021) as:
$\hat{\mathcal{E}}_{\mathcal{D}}^{\gamma}(\theta):=\max_{\|\delta\|\leq\gamma}\frac{1}{N}\sum_{i=1}^{N}\big[L(f(x_{i},\theta+\delta),y_{i})\big].$
@@
$\hat{\mathscr{E}}_{\mathcal{D}}^{\gamma}(\theta):=\max_{\|\delta\|\leq\gamma}\frac{1}{N}\sum_{i=1}^{N}\big[L(f(x_{i}+\delta,\theta),y_{i})\big].$
@@
We now list the necessary propositions from (SWAD:2021) for the rest of the proof.
Proposition 1. For all $\theta\in\Theta,\ |\mathcal{E}_{\mathcal{D}}(\theta)-\mathcal{E}_{\mathcal{T}}(\theta)|\leq\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T}).$
@@
[…]
@@
$\leq\hat{\mathscr{E}}_{\mathcal{D}}^{\gamma_{\mathcal{A}}}(\tilde{\theta}^{*})+\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})+\mathcal{B}(M,N,\delta)$ Def. of $\mathscr{E},\gamma_{\Theta},$ Thm. 1
$\leq\hat{\mathcal{E}}_{\tilde{\mathcal{D}}}(\tilde{\theta}^{*})+\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})+\mathcal{B}(M,N,\delta)$ Def. of $\gamma_{\mathcal{A}},\,\tilde{\theta}^{*}$
C. Experimental Details and Additional Experiments
In this section, we provide a comprehensive account of the experiments corresponding to the tables and figures presented in the main paper, along with supplementary results and discussions that were excluded from the primary manuscript due to page constraints. The additional results include extra robustness evaluations on different backbone networks, clean accuracies of models in Table 3 and 6, and singular value bounds on Theorem 1.
C.1. Experimental Methodology
For all the experiments assessing flatness, mean accuracy, mean corruption error (mCE), and adversarial robustness, we trained three independent models per benchmark and reported the averaged results, with the exception of the ImageNet experiment in Table 5 and 6, which was conducted only once due to high computational demands.
C.2. Flatness Metrics
Descriptions on Metrics. $\mu_{\text{PAC-Bayes}}$ represents simplified PAC-Bayesian bound introduced by PACBayes. The simplified PAC-Bayesian bound, denoted as $\mu_{\text{PAC-Bayes}}$, is computed as $1/\sigma$, where $\sigma$ is the largest value such that $|\mathcal{E}_{\mathcal{D}}(\theta+N(0,\sigma^{2}I))-\mathcal{E}_{\mathcal{D}}(\theta)|\leq\tau$. Therefore, flatter local minima correspond to larger $\sigma$ values and smaller $\mu_{\text{PAC-Bayes}}$. The LPF-based flatness measure evaluates the flatness of a local minimum $\theta^{*}$ by computing the convolution of the loss function $\mathcal{L}$ with a Gaussian kernel $K=N(0,\sigma^{2}I)$. Formally, it is defined as $(\mathcal{L}*K)(\theta^{*})=\int\mathcal{L}(\theta^{*}-\tau)K(\tau)d\tau$. This measure effectively averages the loss over a neighborhood around $\theta^{*}$, with the Gaussian kernel weighting nearby points more heavily, thereby quantifying how sharp the loss landscape is in that region. $\epsilon$-sharpness ($\epsilon_{\text{sharp}}$) measures the sensitivity of a loss function near a local minimum by quantifying the largest perturbation of model parameters that leads to an increase in loss of more than $\epsilon$. We have adopted the default parameters for the $\mu_{\text{PAC-Bayes}}$, LPF measure, and $\epsilon_{\text{sharp}}$ in CIFAR experiments i.e. $\tau=0.05$, $\sigma=0.01$, and $\epsilon=0.1$ respectively. For the larger (tiny)ImageNet runs, we related parameters $\tau=0.35$ and $\epsilon=1.0$ to maintain numerical stability, making the constraints looser than those used for CIFAR.
Training Details. Our training procedure on CIFAR experiments closely follows AugMix9:2021, with the only difference being that we train the WideResNet-40-2 model for 200 epochs instead of 100. We employ an SGD optimizer with an initial learning rate of 0.1, a weight decay of 0.0005, and a momentum of 0.9. Additionally, we use a cosine learning rate decay scheduler that starts with the learning rate of 0.1 and a weight decay of 0.0005. Finally, we evaluate and report the flatness of the converged models. For tinyImageNet and ImageNet experiments, we have majorly adopted the configurations from (AugMix9:2021), with changes in epochs and backbone networks for faster training. We have trained ResNet18 on tinyImageNet and ImageNet for 50 epochs, instead of training ResNet50 for 90 epochs. SGD optimizer with initial learning rate 0.01, weight decay 0.0005, and momentum 0.9 has been adopted. All the models have been trained until convergence, with cosine learning rate decaying scheduling. Regarding the augmentation techniques, default configuration parameters were used for all baseline augmentations except StyleAug, which had hyperparameter $\alpha$ set to $0.95$ for the CIFAR experiments, and 1.0 for the tinyImageNet experiments.
C.3. Robustness Benchmark Details
@@
Adversarial Robustness Benchmarks. To benchmark model robustness under adversarial threat, we evaluated the models trained with augmentations on CIFAR-10/100, tinyImageNet, and ImageNet using widely recognized, untargeted Projected Gradient Descent (PGD) attacks crafted under both $L_{2}$ and $L_{\infty}$ norms. The training environment holds the same as in the common corruption benchmarks. We will use $\epsilon$ and $\alpha$ to denote the maximum allowed perturbation size. Regarding CIFAR-10/100 experiments, we utilized attack configurations from (boundary_thickness). Specifically, we deployed PGD-20 $L_{2}$ attacks with $\epsilon=0.5,\alpha=0.0125$ and PGD-7 $L_{\infty}$ attacks with $\epsilon=8/255,\alpha=2/255$. For tinyImageNet experiments, we utilized PGD-10 $L_{2}$ attacks with $\epsilon=0.25,\alpha=0.025$ and PGD-5 $L_{\infty}$ attacks with $\epsilon=8/255,\alpha=2/255$. On ImageNet experiments, we deployed PGD-10 $L_{2}$ attacks with $\epsilon=0.25,\alpha=0.025$ and PGD-2 $L_{\infty}$ attacks with $\epsilon=2/255,\alpha=1/255$. In essence, milder attack configurations have been deployed for tinyImageNet and ImageNet.
C.4. Additional Experiments on Varying Backbone Networks
To confirm our theorems across different backbone architectures, we provide supplementary experiments on CIFAR-100 robustness benchmarks, expanding upon the data shown in Table 3. On CIFAR-100, augmentations like AugMix, RandAug, and PixMix closely satisfy the PSA condition by producing significant augmented samples in the neighborhood of the original data, which results in flatter minima and improved robustness. In contrast, DeepAug and StyleAug lack augmented samples near the original inputs, failing to satisfy both the PSA condition and improved robustness. This pattern remains consistent across different backbone networks. As illustrated in Table 9, our conclusions remain consistent on distinct backbone architectures (allconvnet; densenet; resnext).
Backbone Benchmark ERM    AugMix    RandAug    PixMix    DeepAug    StyleAug AllConvNet CIFAR-100-C $\downarrow$ 56.88 43.13 (-13.75) 47.52 (-9.36) 37.70 (-19.18) 62.00 (+5.12) 56.90 (+0.02) CIFAR-100-$\overline{\text{C}}$ $\downarrow$ 20.33 9.84 (-10.49) 13.86 (-6.47) 14.29 (-6.04) 64.24 (+43.91) 38.64 (+18.31) CIFAR-100, $L_{2}$ $\downarrow$ 93.96 86.51 (-7.45) 88.96 (-5.00) 74.93 (-19.03) 95.29 (+1.33) 96.56 (+2.60) CIFAR-100, $L_{\infty}$ $\downarrow$ 99.95 99.90 (-0.05) 99.94 (-0.01) 99.64 (-0.31) 99.99 (+0.04) 99.71 (-0.24) DenseNet CIFAR-100-C $\downarrow$ 61.21 43.32 (-17.89) 48.54 (-12.67) 37.58 (-23.63) 61.47 (+0.26) 68.97 (+7.76) CIFAR-100-$\overline{\text{C}}$ $\downarrow$ 23.10 8.87 (-14.23) 13.68 (-9.42) 14.27 (-8.83) 66.20 (+43.10) 43.90 (+20.80) CIFAR-100, $L_{2}$ $\downarrow$ 99.91 97.88 (-2.03) 99.44 (-0.47) 93.19 (-6.72) 98.70 (-1.21) 99.76 (-0.15) CIFAR-100, $L_{\infty}$ $\downarrow$ 100.00 99.99 (-0.01) 99.99 (-0.01) 99.99 (-0.01) 99.99 (-0.01) 99.99 (-0.01) ResNext29 CIFAR-100-C $\downarrow$ 58.05 39.60 (-18.45) 43.49 (-14.56) 31.35 (-26.70) 61.50 (+3.45) 70.11 (+12.06) CIFAR-100-$\overline{\text{C}}$ $\downarrow$ 13.09 3.85 (-9.24) 6.99 (-6.10) 2.64 (-10.45) 66.61 (+53.52) 22.78 (+9.69) CIFAR-100, $L_{2}$ $\downarrow$ 98.72 91.68 (-7.04) 95.95 (-2.77) 91.64 (-7.08) 98.36 (-0.36) 99.67 (+0.95) CIFAR-100, $L_{\infty}$ $\downarrow$ 100.00 99.94 (-0.06) 99.97 (-0.03) 99.90 (-0.10) 99.96 (-0.04) 99.98 (-0.02)
C.5. Results on Theorem 1’s Bound